OpenStudy (crashonce):

Starting from the third term of the sequence 1, 3, 2, ..., each term is equal to the difference of that the term preceding it minus the term preceding that. Find the sum of the first 100 terms of the sequence

OpenStudy (crashonce):

@perl

OpenStudy (perl):

a1 = 1 a2 = 3 a(n) = a(n-1) - a(n-2)

OpenStudy (crashonce):

then?

OpenStudy (crashonce):

ok whats the answer lol

OpenStudy (perl):

a(3) = a(2) - a(1) a(4) = a(3) - a(2) a(5) = a(4) - a(3) do you see how they cancel if you add a(3) + a(4) + a(5)

OpenStudy (crashonce):

ye

Miracrown (miracrown):